# Rollbaccine: Herd Immunity against Storage Rollback Attacks in TEEs (Technical Report)

### v1

@@ 7. Evaluation
Implementation. We implemented Rollbaccine as a device mapper for Linux kernel 6.8 available here: https://anonymous.4open.science/r/rollbaccine-883E (3,938 LoC). AEAD uses in-kernel AES-GCM; hashing uses HMAC-SHA256. We use in-kernel TCP connections with signed messages for primary-backup communication.
@@
We compare Rollbaccine’s performance against four systems: Unreplicated, DM, Replicated, and Nimble. Unreplicated reads and writes from local (ephemeral) disk without replication. It represents the highest-performing but least durable and secure option. DM adds dm-crypt and dm-integrity for encryption and integrity validation, using the same AES-GCM cipher as Rollbaccine. dm-crypt + dm-integrity provides confidentiality and detection of random data corruptions. Replicated uses the highest-performing durable disk available to Azure VM-based TEEs, a locally 3-way replicated P80 Premium SSD rated for 20,000 IOPS. Both DM and Replicated write integrity metadata to disk (azureStorage, ), which is not sufficient against rollback attacks; the integrity hash could be rolled back along with the data by a motivated attacker.
Nimble (nimble, ) is a state-of-the-art solution against rollback attacks that is general, resistant, but not automatic. Applications must be manually modified to send state updates to a “coordinator” that persists the updates to untrusted storage, replicates to 3 TEE-based “endorsers”, and then replies to the application. These modifications are labor-intensive; it took three person-months to modify HDFS into NimbleHDFS (nimble, ).
We evaluate against four configurations of NimbleHDFS: NimbleHDFS-100, NimbleHDFS-100-Mem, NimbleHDFS-1, and NimbleHDFS-1-Mem. The number (100 or 1) represents batch size. The original paper batches and replicating every 100 writes, creating a window of vulnerability during which writes marked “durable” may be rolled back by an attacker (nimble, ), breaking the semantic guarantees of HDFS. Setting batch size to 1 preserves semantics. The -Mem modifier indicates whether state updates are persisted to locally replicated Standard LRS storage as described in the paper or kept in the coordinator’s memory (and not fault tolerant). We co-locate the coordinator machine with the NimbleHDFS to reduce network latency.
Rollbaccine is evaluated with six configurations. Rollbaccine is the standard setup, with $f=1$ and $L=0$ (all 2.4GB of integrity metadata in memory). Rollbaccine-sync synchronously replicates all writes regardless of persistence flags in order to isolate the effect of asynchronous replication. Rollbaccine-f=0 and Rollbaccine-f=2 toggle between no backups (only rollback detecting) and 2 backups, measuring the overhead of networking. Rollbaccine-L=1 and Rollbaccine-L=2 place the bottommost $L$ layers of the integrity metadata Merkle tree on disk, measuring the overhead of read/write amplification, requiring only 0.15GB and 9.6MB of memory for integrity metadata respectively.
@@ 7.1. Performance overview
We evaluate Rollbaccine with the following benchmarks and unmodified applications: (1) TPC-C (tpc-c, ) over PostgreSQL mounted on ext4 (Figure 4), (2) NNThroughputBenchmark (nnThroughputBenchmark, ) over HDFS (hdfs, ) mounted on ext4 (Figure 5(a)), and (3) Filebench (filebench, ) Varmail and Webserver workloads over ext4 and xfs (Figures 5(b) and 5(c)).
Figure 4. PostgreSQL TPC-C throughput-latency graph.
PostgreSQL. PostgreSQL is a widely-used open-source transactional database that guarantees the durability of committed transactions by persisting writes to disk. Rollback attacks on disk can break durability, allowing attackers to remove unwanted transactions. PostgreSQL contains 1.3M LoC, making it infeasible to manually rewrite for rollback resistance. It is therefore a prime target for Rollbaccine, which promises automatic rollback resistance. We benchmark PostgreSQL with TPC-C, a standard OLTP benchmark for transactional databases. We configure TPC-C to run with 10 warehouses and set the isolation level to TRANSACTION_SERIALIZABLE. The results are in Figure 4.
Compared to Unreplicated, DM introduces negligible overhead. Replicated and Rollbaccine respectively reduce throughput by $35\%$ and $15\%$ and increase latency by $54\%$ and $19\%$. This can be attributed to the fact that when benchmarked with TPC-C, roughly every 1 in 5 operations in PostgreSQL are persisted, because every transaction must be durably flushed to PostgreSQL’s Write Ahead Log (WAL) before commit. Both Replicated and Rollbaccine must then synchronously replicate over the network, introducing additional latency overhead, although the latency for Replicated is an order of magnitude greater (§ 7.3). Despite this, the performance penalty is not severe because, at 10 warehouses, TPC-C is contention bottlenecked.
Of the configurations of Rollbaccine, Rollbaccine-sync performs the worst, unable to leverage the benefits of asynchronous replication. The differences between Rollbaccine-f=0, Rollbaccine (with f=1), and Rollbaccine-f=2 illustrate the overhead of networking, whereas the differences between Rollbaccine, Rollbaccine-L=1, and Rollbaccine-L=2 demonstrate the effect of read/write amplification from accessing Merkle tree integrity metadata on disk.
@@
Figure 5. Performance results. Nimble’s configurations contain $\circ$”.
HDFS. Hadoop Distributed File System is the file system backing Hadoop MapReduce. Rollback attacks can break the persistence guarantees of HDFS (hdfs, ). We configure HDFS to run with one namenode and evaluate it with Hadoop’s NNThroughputBenchmark (nnThroughputBenchmark, ); each operation uses 500,000 files (or directories for mkdirs) and 16 client threads (nimble, ). Results are in Figure 5(a).
DM and Rollbaccine perform similarly to Unreplicated, reducing throughput by at most $5\%$ and at times outperforming Unreplicated (attributed to experimental noise). This is because NNThroughputBenchmark, regardless of the number of client threads, uses a single thread to communicate with HDFS in order to isolate the overhead of RPC calls (nnThroughputBenchmark, ). Once enough client threads are launched (16 is enough) on Unreplicated, DM, or Rollbaccine, this single thread becomes the bottleneck, not HDFS.
@@
The throughput and latency trends are similar for ext4 and xfs, so we will discuss them together. We first examine Varmail. Unlike TPC-C (contention bound with 10 warehouses) and NNThroughputBenchmark (bottlenecked on a single thread), Varmail is bottlenecked on disk, so DM, Replicated, and Rollbaccine all experience throughput and (inversely proportional) latency degradations due to the high volume of synchronous writes. Replicated has the highest average latency per operation due to its high fsync latency. Rollbaccine has the second-highest latency, because it must similarly wait for a network round trip, reducing throughput by $71\%$ and increasing latency by 2.7×2.7$\times$2.7 ×. DM does not perform networking but still suffers from synchronously flushing journal entries to disk.
@@ 7.2. Comparison against Nimble
Rollbaccine provides rollback resistance for any program mounted over Rollbaccine’s device mapper. If an application relies on a cloud service for persistence, then the service itself must be mounted over Rollbaccine, requiring buy-in from cloud providers.
Nimble does not require cloud provider buy-in and instead detects rollbacks in existing services. Unlike block devices, cloud services have diverse APIs, so Nimble must sacrifice automation for compatibility. In addition, the amount of integrity metadata that must be preallocated for cloud services is often opaque to the end-user, so Nimble opts to maintain a log of updates instead. This serves Nimble well for the applications it targets (such as Intel HiBench (hibench, )), which can tolerate windows of vulnerability during which data may be rolled back without detection, allowing Nimble to batch log updates.
Unfortunately, the overheads of sequential log replication resurface when batching is disabled for safety. In NNThroughputBenchmark’s write operations (create, mkdirs, open, delete, fileStatus, rename), Rollbaccine outperforms NimbleHDFS-1 by 208×208$\times$208 × (Figure 5(a)).
Azure storage plays a role; both NimbleHDFS-100 and NimbleHDFS-1 underperform their in-memory counterparts by 25×2-5$\times$2 - 5 ×.
Nimble’s performance penalties, however, mainly stem from synchronous, sequential log replication. NimbleHDFS operations that use multiple threads to save files in parallel must still sequentially append, sign, and replicate each log entry. Its use of asymmetric ECDSA-SHA256 signatures allows the log to be publicly verifiable but introduces additional overhead in the critical section. NimbleHDFS’s throughput then becomes a function of its batch size, reducing NimbleHDFS-1’s throughput to double-digits.
@@ 7.3. Microbenchmarks
Figure 6. Throughput-latency graphs of microbenchmarks. Latency is log-scale.
We analyze Rollbaccine’s performance with fio, varying the following parameters: I/O direction (read or write), sequentiality (sequential or random), buffering (O_DIRECT or not), and persistence (synchronous or asynchronous writes). All operations are of size 4K with iodepth 1. We gradually increase the number of fio threads until throughput saturates for each configuration. For each test, we perform 30 seconds of warmup (in order to fill the page cache for buffered workloads), then record statistics for 60 seconds.
Figure 6 displays the throughput (thousands of IOPS) and average completion latency (ms) for each experiment. Each plot point in the graph is annotated with the number of threads used. Note that latency is log scale, so peaks in throughput may appear diminished in some graphs, and that the throughput and latency scales change for each graph.
@@
Direct I/O or persisted writes. When either O_DIRECT or fsync are used for writes, latency increases to the sub-millisecond range and throughput caps at around 75-150,000 IOPS across all tests (Figures 6(b), 6(d), 6(f), 6(h), 6(j) and 6(l)). This is because the disk cannot coalesce writes, either because it receives each operation individually (O_DIRECT) or requires immediate persistence (fsync).
Random access. Random accesses cap out at 50-80,000 IOPS and sub-millisecond latency (Figures 6(a), 6(b), 6(f), 6(j) and 6(i)), with the exception of buffered writes in Figure 6(e). For random buffered reads, the cap is imposed because page prefetching is ineffective for random accesses and each read must be individually serviced by disk. For random buffered writes, latency is an order of magnitude lower, because although the writes are random, they can still be batched in the page cache and immediately returned. Throughput, however, is quickly capped once writes fill the page cache and the disk becomes the bottleneck.
@@
Unreplicated. Reads reach a peak throughput of around 75,000 IOPS and sub-millisecond latency (Figures 6(a), 6(b) and 6(d)), except for buffered sequential reads, which reach 500,000 IOPS and $10^{-2}$ms latency (Figure 6(c)). This is because, unlike the other read workloads, buffered sequential reads can consistently read prefetched pages from the page cache. However, as number of threads increase, each thread (sequentially) reads from a different location on disk, lowering the efficacy of prefetching and capping throughput. This behavior is universal across configurations.
The throughput and latency of Unreplicated is identical for all write workloads with O_DIRECT (Figures 6(f), 6(h), 6(j) and 6(l)), regardless of sequentiality or persistence, since those writes are disk I/O bottlenecked. For buffered, persisted writes (Figures 6(i) and 6(k)), fsync latency spikes and cripples throughput due to the constant flushing of the page cache.
@@
DM. The majority of overhead for DM comes from dm-integrity (dmIntegrity, ), which maintains a journal of write blocks and their integrity metadata on disk. The journal entry is flushed to disk when persistence is required, and data is asynchronously copied from the entry to their actual locations on disk. When a read is requested, if the metadata is not in memory, it must also be fetched from disk.
@@
Finally, we must explain how DM’s throughput continues rising for sequential, persisted writes (Figures 6(k) and 6(l)). This can again be attributed to journaling. Although journal entries must be flushed to disk after an fsync, a single journal entry’s flush can account for the persistence of multiple writes, in effect batching the fsyncs.
@@
Rollbaccine. Reads in Rollbaccine perform similarly to Unreplicated (Figures 6(a), 6(b), 6(c) and 6(d)) because they do not leave the primary, with a maximum of $16\%$ and $21\%$ additional latency and throughput overheads at saturation, respectively. This overhead is the result of decryption and maintaining the list of invoked and pending operations; the latter happens in a critical section (§ 6.1).
For asynchronous writes, Rollbaccine scales with the number of threads alongside Unreplicated, with a maximum latency and throughput overhead of $43\%$ and $45\%$ respectively (Figures 6(e), 6(f), 6(g) and 6(h)). With the exception of sequential buffered writes, which is bottlenecked on bandwidth (Figure 6(g)), the primary’s disk is the bottleneck. These results demonstrate that by replicating asynchronous writes in the background, Rollbaccine is able to scale.
Persisted writes, on the other hand, are bottlenecked on round-trip time to the backups, with a maximum of $433\%$ and $45\%$ latency and throughput overhead (Figures 6(i), 6(j), 6(k) and 6(l)). Latency increases by an order of magnitude as the primary waits for the backup to receive all previous operations before acknowledging the write. Throughput, however, can continue to scale due to this optimization: if multiple synchronous writes concurrently arrive at the backup, then it only acknowledges the write with the highest index, since that acknowledgment implies the receipt of all writes with lower indices.
In summary, Rollbaccine adds $21\%$ overhead for reads, $45\%$ overhead for asynchronous writes, similar to DM (with the exception of direct writes), and an order of magnitude of overhead for synchronous writes. The under-performance of direct writes is not fundamental; Rollbaccine can be modified to cache writes in memory as well. For synchronous writes, Rollbaccine experiences much higher overheads, but, as seen in § 7.1, most applications are designed to use persistence operations sparingly and are minimally affected.

### v2

@@ 7. Evaluation
Implementation. We implemented Rollbaccine as a device mapper for Linux kernel 6.8 available at https://github.com/davidchuyaya/rollbaccine (3,938 LoC). AEAD uses in-kernel AES-GCM; hashing uses HMAC-SHA256. We use in-kernel TCP connections with signed messages for primary-backup communication.
@@
We compare Rollbaccine’s performance against four systems: Unreplicated, DM, Replicated, and Nimble. Unreplicated reads and writes from local (ephemeral) disk without replication. It represents the highest-performing but least durable and secure option. dm-crypt + dm-integrity provides confidentiality and detection of random data corruptions. DM adds dm-crypt and dm-integrity for encryption and detection of random data corruptions, using the same AES-GCM cipher as Rollbaccine. Replicated uses the highest-performing durable disk available to Azure VM-based TEEs, a locally 3-way replicated P80 Premium SSD rated for 20,000 IOPS. Both DM and Replicated write integrity metadata to disk (azureStorage), which is not sufficient against rollback attacks; the integrity hash could be rolled back along with the data by a motivated attacker.
Nimble (nimble) is a state-of-the-art solution against rollback attacks that is general, resistant, but not automatic. Applications must be manually modified to send state updates to a “coordinator” that persists the updates to untrusted storage, replicates to 3 TEE-based “endorsers”, and then replies to the application. These modifications are labor-intensive; it took three person-months to modify HDFS into NimbleHDFS (nimble).
We evaluate against four configurations of NimbleHDFS: NimbleHDFS-100, NimbleHDFS-100-Mem, NimbleHDFS-1, and NimbleHDFS-1-Mem. The number (100 or 1) represents batch size. The original paper batches and replicates every 100 writes, creating a window of vulnerability during which writes marked “durable” may be rolled back by an attacker (nimble), breaking the semantic guarantees of HDFS. Setting batch size to 1 preserves semantics. The -Mem modifier indicates whether state updates are persisted to locally replicated Standard LRS storage as described in the paper or kept in the coordinator’s memory (and not fault tolerant). We co-locate the coordinator machine with the NimbleHDFS to reduce network latency.
Rollbaccine is evaluated with seven configurations. Rollbaccine is the standard setup, with $f=1$ and $L=0$ (all 2.4GB of integrity metadata in memory). Rollbaccine-multicloud uses a GCP n2d-standard-16 machine (16 vCPUs, 64GB RAM, AMD SEV-SNP TEE) in the West Europe region as the backup to evaluate the cost of cross-cloud deployments. Cross-cloud deployments guarantee rollback resistance even in the presence of Type II attackers (malicious cloud providers). Its ping time to the Azure machine is 23ms. Rollbaccine-sync synchronously replicates all writes regardless of persistence flags in order to isolate the effect of asynchronous replication. Rollbaccine-f=0 and Rollbaccine-f=2 toggle between no backups (only rollback detecting) and 2 backups, measuring the overhead of networking. Rollbaccine-L=1 and Rollbaccine-L=2 place the bottommost $L$ layers of the integrity metadata Merkle tree on disk, measuring the overhead of read/write amplification, requiring only 0.15GB and 9.6MB of memory for integrity metadata respectively.
@@ 7.1. Performance Overview
We evaluate Rollbaccine with the following benchmarks and unmodified applications: TPC-C (tpc-c) over PostgreSQL mounted on ext4 (Figure 4), NNThroughputBenchmark (nnThroughputBenchmark) over HDFS (hdfs) mounted on ext4 (Figure 5(a)), and Filebench (filebench) Varmail and Webserver workloads over ext4 and xfs (Figures 5(b) and 5(c)).
Figure 4. PostgreSQL TPC-C throughput-latency graph with 20, 30, 40, and 50 clients.
PostgreSQL. PostgreSQL is a transactional database that guarantees the durability of committed transactions by persisting writes to disk. Rollback attacks on disk can break durability, allowing attackers to remove unwanted transactions. PostgreSQL contains 1.3M LoC, making it infeasible to manually rewrite for rollback resistance. It is therefore a prime target for Rollbaccine, which promises automatic rollback resistance. We benchmark PostgreSQL using TPC-C with 10 warehouses and isolation level TRANSACTION_SERIALIZABLE. The results are in Figure 4. Latency is log-scale.
Compared to Unreplicated, DM introduces negligible overhead. Replicated and Rollbaccine respectively reduce throughput by $35\%$ and $15\%$ and increase latency by $54\%$ and $19\%$. This can be attributed to the fact that when benchmarked with TPC-C, roughly every 1 in 5 operations in PostgreSQL are persisted, because every transaction must be durably flushed to PostgreSQL’s Write Ahead Log (WAL) before commit. Both Replicated and Rollbaccine must then synchronously replicate over the network, introducing additional delay, although the latency for Replicated is an order of magnitude greater (§ 7.3). Despite this, the performance penalty is not severe because, at 10 warehouses, TPC-C is contention bottlenecked.
Of the configurations of Rollbaccine, Rollbaccine-multicloud has the worst performance due to high inter-cloud latency. Rollbaccine-sync is also an outlier, as it is unable to leverage the benefits of asynchronous replication. This represents a lower bound on performance. The remaining configurations of Rollbaccine take advantage of the observation that applications, such as PostgreSQL, are already designed to minimize persistence and carefully choose when to fsync, so blocking on replication is only necessary for persistent writes.
The differences between Rollbaccine-f=0, Rollbaccine (with f=1), and Rollbaccine-f=2 illustrate the overhead of networking, whereas the differences between Rollbaccine, Rollbaccine-L=1, and Rollbaccine-L=2 demonstrate the effect of read/write amplification from accessing Merkle tree integrity metadata on disk.
@@
Figure 5. Performance results. Each bar is labeled with the exact throughput on top.
HDFS. Hadoop Distributed File System is the file system backing Hadoop MapReduce. Rollback attacks can break the persistence guarantees of HDFS (hdfs). We configure HDFS to run with one namenode and evaluate it with Hadoop’s NNThroughputBenchmark (nnThroughputBenchmark); each operation uses 500,000 files (or directories for mkdirs) and 16 client threads (nimble). Results are in Figure 5(a).
DM and Rollbaccine perform similarly to Unreplicated, reducing throughput by at most $5\%$ and at times outperforming Unreplicated (attributed to experimental noise). This is because NNThroughputBenchmark, regardless of the number of client threads, uses a single thread to communicate with HDFS in order to isolate the overhead of RPC calls (nnThroughputBenchmark). Once enough client threads are launched (16 is enough) on Unreplicated, DM, or Rollbaccine, this single thread becomes the bottleneck, not HDFS.
@@
The throughput and latency trends are similar for ext4 and xfs, so we will discuss them together. We first examine Varmail. Unlike TPC-C (contention bound with 10 warehouses) and NNThroughputBenchmark (bottlenecked on a single thread), Varmail is bottlenecked on disk, so DM, Replicated, and Rollbaccine all experience throughput and (inversely proportional) latency degradations due to the high volume of synchronous writes. Replicated has the highest average latency per operation due to its high fsync latency. Rollbaccine has the second-highest latency, because it must similarly wait for a network round trip, reducing throughput by $71\%$ and increasing latency by $2.7\times$. DM does not perform networking but still suffers from synchronously flushing journal entries to disk.
@@ 7.2. Comparison against Nimble
Nimble implements rollback resistance by maintaining a replicated log; applications must be modified in order to append state updates to the log. Nimble then batches those updates in order to improve performance; the way it batches, however, is incompatible with block device crash consistency. If the batch size is set to $B$, then replication occurs every $B$ writes; the first B1B-1 writes of each batch, persistent or not, will be returned immediately before they are made rollback resistant. NimbleHDFS-1, with a batch size of 1, is therefore the only configuration that is safe.
We ask whether a general purpose rollback tolerant solution like Rollbaccine can match the performance of a manual, rollback-resistant solution. We compare against NimbleHDFS-1, the only version of Nimble to provide block device crash consistency. Our results were surprising. Rollbaccine not only matches but outperforms NimbleHDFS-1 by $208\times$ on NNThroughputBenchmark’s write operations (create, mkdirs, delete, rename) (Figure 5(a)). The discrepancy holds even when comparing against Rollbaccine-sync, where we have Rollbaccine artificially replicate every disk write synchronously.
Azure storage plays a role; both NimbleHDFS-100 and NimbleHDFS-1 underperform their in-memory counterparts by $2-5\times$.
But the culprit is CPU (§ 7.1). Round-trip network latency introduces negligible overhead on the CPU-bottlenecked thread (which is why Rollbaccine and Rollbaccine-sync perform similarly), whereas the asymmetric ECDSA-SHA256 signatures used by Nimble’s messages overwhelm it. NimbleHDFS’s throughput then becomes a function of how many messages it must send; larger batch sizes allow it to amortize signatures across log entries, but once batching is disabled for safety, its throughput reduces to double-digits.
@@ 7.3. Microbenchmarks
Figure 6. Low contention throughput-latency graphs annotated with the number of threads. Latency is log-scale.
We analyze Rollbaccine’s performance with fio, varying I/O direction (read or write), sequentiality (sequential or random), buffering (O_DIRECT or not), persistence (synchronous or asynchronous writes), and contention (whether all threads write to the same part of the device or not). All operations are of size 4K with iodepth 1. We gradually increase the number of fio threads until throughput saturates for each configuration. For each test, we perform 30 seconds of warmup (filling the page cache), then record statistics for 60 seconds.
Figures 6 and 7 display the throughput (thousands of IOPS) and average completion latency (ms) of low- and high-contention experiments, respectively. Each plot point in the graph is annotated with the number of threads used. Note that latency (and throughput in Figure 7) is log scale, and that the throughput and latency scales change for each graph.
@@
Direct I/O or persisted writes. When either O_DIRECT or fsync are used for writes, latency increases to the sub-millisecond range and throughput caps at around 75-150,000 IOPS across all tests (Figures 6(b), 6(d), 6(f), 6(h), 6(j), 6(l) and 7) except for high contention reads over DM, explained below. This is because the disk cannot coalesce writes, either because it receives each operation individually (O_DIRECT) or requires immediate persistence (fsync).
Random access. Random accesses cap out at 50-80,000 IOPS and sub-millisecond latency (Figures 6(a), 6(b), 6(f), 6(j) and 6(i)), with the exception of buffered writes in Figure 6(e). For buffered reads, the cap is imposed because page prefetching is ineffective for random accesses and each read must be individually serviced by disk. For buffered writes, latency is an order of magnitude lower, because although the writes are random, they can still be batched in the page cache and immediately returned. Throughput, however, is quickly capped once writes fill the page cache and the disk becomes the bottleneck.
((a)) Write, direct
((b)) Read, direct
Figure 7. High-contention graphs similar to Figure 6. Both latency and throughput are log-scale.
@@
Unreplicated. Reads reach a peak throughput of around 75,000 IOPS and sub-millisecond latency (Figures 6(a), 6(b) and 6(d)), except for buffered sequential reads, which reach 500,000 IOPS and $10^{-2}$ms latency (Figure 6(c)), and high contention reads, which are capped at 18,000 IOPS (Figure 7(b)). Buffered sequential reads benefit from prefetching, although as the number of threads increase, each thread (sequentially) reads from a different location on disk, lowering the efficacy of prefetching and capping throughput. This behavior is universal across configurations. High contention reads fail to scale across all configurations (except for DM, explained below) due to read collisions at the SSD level (ssdPerformance; ssdReadCollisions).
The throughput and latency of Unreplicated is identical for all write workloads with O_DIRECT (Figures 6(f), 6(h), 6(j), 6(l) and 7(a)), regardless of sequentiality, persistence, or contention, since those writes are disk I/O bottlenecked. For buffered, persisted writes (Figures 6(i) and 6(k)), fsync latency spikes and cripples throughput due to the constant flushing of the page cache.
@@
DM. The majority of overhead for DM comes from dm-integrity (dmIntegrity), which maintains a journal of write blocks and their integrity metadata on disk. The journal entry is flushed to disk when persistence is required, and data is asynchronously copied from the entry to their actual locations on disk. When a read is requested, if the metadata is not in memory, it must also be fetched from disk.
@@
For sequential, persisted writes (Figures 6(k) and 6(l)), DM’s throughput continues rising due to its journaling. Although journal entries must be flushed to disk after an fsync, a single journal entry’s flush can account for the persistence of multiple writes, in effect batching the fsyncs.
Journaling also allows DM to minimize disk accesses when it comes to high contention (Figure 7). Reads from journal entries already in-memory can be serviced without going to disk, and writes can be returned immediately after a journal entry is created in-memory.
@@
Rollbaccine. Reads in Rollbaccine perform similarly to Unreplicated (Figures 6(a), 6(b), 6(c) and 6(d)) because they do not leave the primary, with a maximum of $16\%$ and $21\%$ additional latency and throughput overheads as the result of decryption and maintaining the list of invoked and pending operations; the latter happens in a critical section (§ 6.1). High contention reads (and writes) are the exception (Figure 7), as the conflicting operations are sequentially executed.
For asynchronous, low contention writes, Rollbaccine scales with the number of threads alongside Unreplicated, with a maximum latency and throughput overhead of $43\%$ and $45\%$ respectively (Figures 6(e), 6(f), 6(g) and 6(h)). With the exception of sequential buffered writes, which is bottlenecked on bandwidth (Figure 6(g)), the primary’s disk is the bottleneck. These results demonstrate that by replicating asynchronous writes in the background, Rollbaccine is able to scale.
Persisted writes, on the other hand, are bottlenecked on round-trip time to the backups, with a maximum of $433\%$ and $45\%$ latency and throughput overhead (Figures 6(i), 6(j), 6(k) and 6(l)). Latency increases by an order of magnitude as the primary waits for the backup to receive all previous operations before acknowledging the write. Throughput, however, can continue to scale due to this optimization: if multiple synchronous writes concurrently arrive at the backup, then it only acknowledges the write with the highest index, since that acknowledgment implies the receipt of all prior writes.
In summary, Rollbaccine adds $21\%$ overhead for reads, $45\%$ overhead for asynchronous writes, similar to DM (with the exception of direct writes), and an order of magnitude of overhead for synchronous writes and high contention operations. The under-performance of direct writes and high contention operations is not fundamental; Rollbaccine can be modified to cache and service reads and writes from memory similar to DM. For synchronous writes, Rollbaccine experiences much higher overheads, but, as seen in § 7.1, most applications are designed to use persistence operations sparingly and are minimally affected.